# Cohomology of the pure symmetric automorphisms of right-angled Artin groups

### 15 Apr 2026 v1

@@ 5. Computing $H^{*}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$
For free abelian groups, $H^{*}\left(\mathbb{Z}^{n}\right)$ is an exterior algebra generated by one-dimensional classes, so the set $\left\{\bar{\gamma}_{I}^{j}\mid\bar{C}_{I}^{j}\in\mathcal{B}(\tau)\right\}$ is a generating set for the cohomology of $\mathrm{Stab}(\sigma)$. As a consequence, the set of all products of $q$ distinct $\bar{\gamma}_{I}^{j}$ classes is a generating set for $H^{q}(\operatorname{Stab}(\sigma))$. Therefore, we may choose as generating set for $H^{q}(\operatorname{Stab}(\sigma))$ the collection of dual clases of all subsets of canonical outer-automorphisms $\mathcal{A}\subset\mathcal{B}(\tau)$ with $|\mathcal{A}|=q$. Taking all of this into account, we see that in $E_{1}^{p,q}$ we have a basis element for each compatible collection $\mathcal{A}$ of canonical outer-automorphisms with $|\mathcal{A}|=q$ and each $\sigma$ of dimension $p$ where $\mathcal{A}\subseteq\mathcal{B}(\sigma)$.
@@ 5.1. The essential filtration
rangin over the set of compatible families $\mathcal{A}$, with $|\mathcal{A}|=q$ and $s(\mathcal{A})=s$; and
@@ 5.2. The $E_{2}$ page
The algebra $H^{*}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$ is generated by the one-dimensional classes $\bar{\gamma}_{I}^{j}$ where the $\bar{C}_{I}^{j}$ are canonical outer-automorphisms.
@@
Observe that even if $A_{\Gamma}$} looks close to $F_{4}$, the differences in the homology of the symmetric oouter-automorphisms group are remarkable: by the formula given in Lemma 5.2 in [9], we get
@@ 6.2. Applying the Leray-Hirsch Theorem
$E_{2}^{*,q}=H^{*}\left(\Sigma\mathrm{POut}(A_{\Gamma}),H^{q}\left(A_{\Gamma}\right)\right)=H^{*}\left(\Sigma\mathrm{POut}(A_{\Gamma}),\mathbb{Z}^{t_{q}}\right)$
@@
$\iota^{*}:H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\rightarrow H^{*}\left(A_{\Gamma}\right)$
@@
$H^{1}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\rightarrow H^{1}\left({A_{\Gamma}}\right)$
@@
$H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\rightarrow H^{*}\left({A_{\Gamma}}\right)$
@@
Let $K_{i}$ be the number of k essential vertex types of rank $i$ (generators of $H^{i}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$) and $N_{j}$ the number of $j$-cliques in $\Gamma$ (generators of $H^{j}\left({A_{\Gamma}}\right)$, with $N_{0}=1$). Then
$H^{q}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\cong\bigoplus_{i+j=q}H^{i}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)\otimes H^{j}\left({A_{\Gamma}}\right)$
$\cong\bigoplus_{i+j=q}H^{i}\left(\Sigma\mathrm{POut}(A_{\Gamma}),H^{j}\left(A_{\Gamma}\right)\right)\cong\bigoplus_{i+j=q}\mathbb{Z}^{K_{i}\cdot N_{j}}$
@@
The ring $H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$ is generated in degree 1.
@@ 7. The cohomology ring $H^{*}(\Sigma\mathrm{PAut}({A_{\Gamma}}))$
With the previous notation, relation $(1)$ holds in $H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$.
@@
$\overline{\varphi}_{1}^{*}:H^{1}\left(F_{2}\right)\rightarrow H^{1}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right).$
@@
With the notation established before, relation $(2)$ also holds in $H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$.
@@
$\phi:H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\twoheadrightarrow R.$
@@
(2) $H^{2}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\cong\bigoplus_{i+j=2}\mathbb{Z}^{K_{i}\cdot N_{j}}\cong\mathbb{Z}^{K_{2}\cdot N_{0}}\oplus\mathbb{Z}^{K_{1}\cdot N_{1}}\oplus\mathbb{Z}^{K_{0}\cdot N_{2}}$
@@
Conjecture 7.10 (Generalized Browstein-Lee Conjecture).
@@ 8. The Torelli subgroup
$H^{*}\left(\mathrm{IA}({A_{\Gamma}})\right)\twoheadrightarrow H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right).$

### 16 Apr 2026 v2

@@ 5. Computing $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$
For free abelian groups, $H^{*}(\mathbb{Z}^{n})$ is an exterior algebra generated by one-dimensional classes, so the set $\left\{\bar{\gamma}_{I}^{j}\mid\bar{C}_{I}^{j}\in\mathcal{B}(\tau)\right\}$ is a generating set for the cohomology of $\mathrm{Stab}(\sigma)$. As a consequence, the set of all products of $q$ distinct $\bar{\gamma}_{I}^{j}$ classes is a generating set for $H^{q}(\operatorname{Stab}(\sigma))$. Therefore, we may choose as generating set for $H^{q}(\operatorname{Stab}(\sigma))$ the collection of dual classes of all subsets of canonical outer-automorphisms $\mathcal{A}\subset\mathcal{B}(\tau)$ with $|\mathcal{A}|=q$. Taking all of this into account, we see that in $E_{1}^{p,q}$ we have a basis element for each compatible collection $\mathcal{A}$ of canonical outer-automorphisms with $|\mathcal{A}|=q$ and each $\sigma$ of dimension $p$ where $\mathcal{A}\subseteq\mathcal{B}(\sigma)$.
@@ 5.1. The essential filtration
ranging over the set of compatible families $\mathcal{A}$, with $|\mathcal{A}|=q$ and $s(\mathcal{A})=s$; and
@@ 5.2. The $E_{2}$ page
The algebra $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$ is generated by the one-dimensional classes $\bar{\gamma}_{I}^{j}$ where the $\bar{C}_{I}^{j}$ are canonical outer-automorphisms.
@@
Observe that even if $A_{\Gamma}$} looks close to $F_{4}$, the differences in the homology of the symmetric outer-automorphisms group are remarkable: by the formula given in Lemma 5.2 in [9], we get
@@ 6.2. Applying the Leray-Hirsch Theorem
$E_{2}^{*,q}=H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}),H^{q}(A_{\Gamma}))=H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}),\mathbb{Z}^{t_{q}})$
@@
$\iota^{*}:H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\rightarrow H^{*}(A_{\Gamma})$
@@
$H^{1}(\Sigma\mathrm{PAut}(A_{\Gamma}))\rightarrow H^{1}({A_{\Gamma}})$
@@
$H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\rightarrow H^{*}({A_{\Gamma}})$
@@
Let $K_{i}$ be the number of k essential vertex types of rank $i$ (generators of $H^{i}(\Sigma\mathrm{POut}(A_{\Gamma}))$) and $N_{j}$ the number of $j$-cliques in $\Gamma$ (generators of $H^{j}({A_{\Gamma}})$, with $N_{0}=1$). Then
$H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))\cong\bigoplus_{i+j=q}H^{i}(\Sigma\mathrm{POut}(A_{\Gamma}))\otimes H^{j}({A_{\Gamma}})$
$\cong\bigoplus_{i+j=q}H^{i}(\Sigma\mathrm{POut}(A_{\Gamma}),H^{j}(A_{\Gamma}))\cong\bigoplus_{i+j=q}\mathbb{Z}^{K_{i}\cdot N_{j}}$
@@
The ring $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ is generated in degree 1.
@@ 7. The cohomology ring $H^{*}(\Sigma\mathrm{PAut}({A_{\Gamma}}))$
With the previous notation, relation $(1)$ holds in $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$.
@@
$\overline{\varphi}_{1}^{*}:H^{1}(F_{2})\rightarrow H^{1}(\Sigma\mathrm{PAut}(A_{\Gamma})).$
@@
With the notation established before, relation $(2)$ also holds in $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$.
@@
$\phi:H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\twoheadrightarrow R.$
@@
(2) $H^{2}(\Sigma\mathrm{PAut}(A_{\Gamma}))\cong\bigoplus_{i+j=2}\mathbb{Z}^{K_{i}\cdot N_{j}}\cong\mathbb{Z}^{K_{2}\cdot N_{0}}\oplus\mathbb{Z}^{K_{1}\cdot N_{1}}\oplus\mathbb{Z}^{K_{0}\cdot N_{2}}$
@@
Conjecture 7.10 (Generalized Brownstein-Lee Conjecture).
@@ 8. The Torelli subgroup
$H^{*}(\mathrm{IA}({A_{\Gamma}}))\twoheadrightarrow H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma})).$